Given 812.5g of Sr2P3, how many atoms of strontium oxide can you produce?
1
Expert's answer
2020-12-16T05:52:01-0500

Sr3P2+2H2O=2Sr(OH)2+2PH3

find n Sr3P2

"n=\frac{m}{M}=\frac{812.5}{268.1620}=3.03"

find n Sr(OH)2


"\frac{3.03}{1}=\frac{x}{2}"


x=6.06

Sr(OH)2=SrO+H2O

find n SrOH

"\frac{6.06}{1}=\frac{x}{1}"

x=6.06

find atoms SrO :

"N=Na\times n=6.022\times10^{23}\times6.06=36.49\times10^{23}"